OpenStudy (anonymous):

HELP !! GIVING OUT MEDALS ;) NEED ANSWER NOW!! Evaluate 2b^2 - 3b + 4a^2 for a = -6 and b =8 A. -40 B. 202 C. 248 D. 346

OpenStudy (anonymous):

You have to insert the a and b in your equation

OpenStudy (anonymous):

yeah just plug in soo 2(8^2) - 3(8) + 4 x-6^2

OpenStudy (anonymous):

ya so when you plug in the values you get 248

OpenStudy (anonymous):

-40

OpenStudy (anonymous):

depends on if you use parenthesis lol

OpenStudy (anonymous):

I might have plugged in wrong

OpenStudy (anonymous):

2*8² -3*8 +4*(-6)² = 2*64 -24 + 4*36 =

OpenStudy (anonymous):

but if you plug in correct you should get one of the answers above,when i plugged in I got -40 so just plug in and do PEMDAS

OpenStudy (anonymous):



OpenStudy (anonymous):

128-24 + 148 = 252 (-6)² = -6*-6 = +36

OpenStudy (anonymous):

exactly, 2*64= 128 3*8=24 4*36=144 so 128-24+144=248

OpenStudy (anonymous):

when negative numbers are squared they become positive

OpenStudy (anonymous):

Is it 248 , or -40?

OpenStudy (anonymous):

sorry my 148 should have been 144 so 248 definitely

OpenStudy (anonymous):

yeah , its 248 thanks